If y varies inversely as 3x +2 and y=24 when x =1 , find x when y is 15?
Help a guy out?

Answer by stanbon(75887)

If y varies inversely as 3x +2 and y=24 when x =1
--------------------------------------
y = k/(3x+2)
------------------------
Solve for "k":
24 = k/(3*1+2)
24 = k/5
k = 120
---
Equation:
y = 120/(3x+2)
-----
Find x when y is 15?
15 = 120/(3x+2)
Cross-multiply:
3x+2 = 120/15
3x + 2 = 8
3x = 6
x = 2
===============================
Cheers,
Stan H.